OpenStudy (anonymous):

can someone help me with this please: lim x-->5 cube root (3x^2-4x+9)

OpenStudy (anonymous):

4

OpenStudy (anonymous):

Since the cube root function is continuous for all x and so is the polynomial radicand, we can find the limit by direct substitution of x=5. 3*5^2-4(5)+9=75-20+9=64 and the cube root of 64 is 4. So your answer is in fact 4.

OpenStudy (anonymous):

\[\sqrt[3]{3*5^2-4*5+9}=\sqrt[3]{75-20+9}=\sqrt[3]{64}=4\]

OpenStudy (lgbasallote):

this is solveable via scientific calculator

OpenStudy (anonymous):

thank you!